Bennoch Corporation is expected to pay $2 dividends per share next year (year 1) to its shareholders. Its required rate of return on equity is 10%. Dividends are expected to grow at 5% per year for year 2 through year 3, and then slow down to a steady long-term growth rate of 2% for year 4 and beyond. What is the fair value of its stock price today?
Answer #1

D1=2

D2=(2*1.05)=2.1

D3=(2.1*1.05)=2.205

Value after year 3=(D3*Growth rate)/(Required return-Growth rate)

=(2.205*1.02)/(0.1-0.02)

=28.11375

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2/1.1+2.1/1.1^2+2.205/1.1^3+28.11375/1.1^3

=$26.33(Approx).